OpenStudy (enricoanders):

Find the equation in standard form of the line with slope 5/7 that passes through the point (10,9) Hint: Remember that A, B, and C must be integers, not fractions

OpenStudy (enricoanders):

@welshfella Help

OpenStudy (welshfella):

use the point slope equation for a sstraight line y - y1 = m(x - x1) here m = 5/7 , x1 = 10 and y1 = 9

OpenStudy (welshfella):

this will give you